# ChartX & ChartVLM: A Versatile Benchmark and Foundation Model for Complicated Chart Reasoning

### v5

@@ 1 Introduction
Versatile Multi-modal Large Language Models (MLLMs) [4, 31, 1, 7, 3, 41, 34] have made promising progress in general-purpose vision-language applications such as multi-modal Question Answering (QA) [21, 4, 39, 25], embodied AI [14], and mathematical reasoning [37, 43, 15]. Although MLLMs have demonstrated their powerful generalization ability in a wide range of multi-modal tasks, their performance in multi-modal reasoning tasks still falls short of human abilities [44, 6, 38, 1]. For instance, humans can easily extract numerical values from a given visual chart and engage in a series of complicated logical reasoning based on the extracted values. However, at present, the MLLMs’ ability to perform complicated logical reasoning based on chart data has not been fully explored.
@@
On top of this, the existing open-source chart datasets are consolidated for the training of ChartVLM, including ChartQA [27], Chart-to-text [30], PlotQA [29], and SimChart9K [40]. Note that during the training process, ChartVLM has no access to any data from the ChartX evaluation set. Then, we conduct comprehensive comparisons of ChartVLM with current MLLMs [4, 25] on the ChartX evaluation set, including base abilities, e.g., data extraction, and advanced abilities, e.g., complicated problem-solving, where we demonstrate the superiority of our ChartVLM.
@@ 2 Related Work
Figure 1: Our work offers two insights: a) ChartX: a comprehensive multi-modal chart evaluation set encompassing 22 disciplinary topics, 18 chart types, and 7 tasks where models are evaluated using task-specific metrics such as EM, GPT-acc GPT-score, SCRM [40], and b) ChartVLM: a novel framework to perform the multi-tasks in the chart domain. Our key point is to leverage the instruction adapter to dynamically choose the task that needs to be executed. For downstream tasks that rely on querying chart information, we prioritize chart structural extraction before engaging in chart reasoning tasks. This sequence aims to enhance the interpretability of the reasoning results.
Table 1: Comparison with the existing chart-related benchmarks, where ChartX is constructed for comprehensively evaluating the off-the-shelf vision-language large models from more chart types and topics. Besides, EM denotes Exact Match and SCRM represents the Structuring Chart-oriented Representation Metric described in StructChart [40].
@@
PlotQA [29] N/A 3 1 33.7K PlotQA EM & AP
Chart-to-text [30] 6 6 1 6.6K Chart-to-text EM
ChartQA [27] 15 3 1 1.5K ChartQA EM
OpenCQA [18] 10 5 1 1.2K OpenCQA EM & BLEU & ROUGE
@@
ChartLlama [8] N/A 10 7 1.5K ChartQA & Chart-to-text EM & GPT
ChartBench [42] N/A 9 4 2K ChartBench Accuracy
MMC [24] 5 6 9 2.1K MMC GPT
ChartAssisstant [28] N/A 9 5 1.5K ChartQA & OpenCQA EM& BLEU
@@
Chart Perception aims to extract the numerical and textual information from a given visual chart. By leveraging the OCR tools [26] to supplement the textual information, the basic function of extracting chart information can be achieved. Recently, some researchers [9, 33, 13] have attempted to perform a chart-to-table transformation for the visual chart perception task, by means of self-supervision from image-table pairs. For example, Deplot [22] fine-tuned an image-to-text transformer for such conversion. StructChart [40] utilizes the encoder-decoder framework to achieve transformation. These methods extract the tabular format of a visual chart and leverage the external module such as GPT [32, 5] to perform downstream tasks. However, their chart-related reasoning abilities strongly depend on external modules, whose scalability is hard to guarantee.
Chart Cognition is defined as a process to deal with intricate tasks related to both chart-related knowledge and common sense knowledge. A typical example is to query numerical points from a chart and give the prediction results using mathematical or logical reasoning. Recent studies [10, 36, 45, 19, 23] focus on showing the reasoning ability of their models on chart domain. Pix2Struct [19] presents a pre-training method using masked screenshots from web pages, which is verified to be effective in chart understanding tasks such as ChartQA dataset [27]. Besides, MatCha [27] decodes the answers to chart questions in an end-to-end manner, where the chart reasoning ability can be enhanced from MATH data [35].
Multi-Modal Chart Generation and Benchmark. Chart data generation is a crucial step for scaling up the model ability [36, 23, 2]. Previous chart-related benchmarks only cover general three types of charts (line, pie, bar charts) and focus on a few tasks such as chart-to-table tasks for ChartQA [27], PlotQA [29], and Chart-to-Text [30], and QA tasks for DVQA [16] and OpenCQA [18]. Recently, various benchmarks have been proposed in some works, e.g. MMC [24], ChartLlama [8], ChartBench [42], and ChartAssisstant [28], with the common characteristics of more types, more tasks, and more modalities of chart data, which is insightful for the chart community. However, as shown in Table 1, the data and metric diversity of charts used for evaluating multi-modal large models is relatively limited. For example, ChartBench [42] merely uses a two-sided judgment (yes or no) to evaluate model performance. The types of charts and data in MMC [24] are also insufficient in verifying the chart ability of the off-the-shelf MLLMs.
@@ 3.1 Coverage Analysis of the Evaluation Set
Chart Tasks. Unlike previous chart benchmarks focusing on the category of visual logic reasoning tasks, the ChartX benchmark emphasizes the interpretability for all downstream chart-related tasks. Given that interpretability relies heavily on the ability to perceive chart information, ChartX categorizes perception-related tasks as base tasks, including title perception, chart type recognition, and Structural Extraction (SE). On the other hand, other chart-related tasks are classified as intricate cognition tasks, including chart-related Question Answering (QA), Chart Description, Chart Summarization, and Chart Redrawing. In the context of ChartX, QA refers to answering questions that are formulated solely based on the chart data, requiring reasoning derived directly from the provided chart information. This characteristic distinguishes ChartX from previous chart-related QA datasets like ChartQA [27]. In ChartQA [27], there exists a certain number of QA pairs that cannot be answered solely based on the information presented in the given chart image. Chart Description aims at presenting detailed information and some insights from the distribution of chart data, while Chart Summarization features summarizing the trend-like or high-level characteristics from the given data in a few sentences. Chart Redrawing refers to plotting the given data into a new chart image with the same chart type of original data. The distribution of each task is listed in Fig. 1. For each image, together with labels of base tasks, we collect two QA samples, one description sample, one summarization sample, and one redrawing code sample. Overall, the samples from multi-tasks reach 48K in ChartX.
@@ 3.3 Two-stage Chart Data Generation
Utilizing the strong generation capabilities of GPT-4 [1], ChartX is created through an automated online generation process with manual instructions. This involves a data-centric two-stage generation paradigm, encompassing the creation of perception and cognition data.
@@ 3.4 Task Evaluation Metrics
SCRM. Given that data in the chart has matrix-like row-column transformation invariance and transpose transformation invariance, Structuring Chart-oriented Representation Metric (SCRM) [40] is employed to evaluate the extracted chart information (i.e. SE task), in which the linearized CSV tokens predicted by all models will be transformed to triplet format for performing SCRM evaluation.
@@ 4.3 Cascaded Decoders Design
For basic perception tasks, we fine-tune all the network weights pre-trained from Pix2Struct-base and Pix2Struct-large [19] model, using image-CSV pair data. The fine-tuned encoder and decoder are regarded as chart image encoder and base decoder in ChartVLM. After the fine-tuning stage is completed, the encoder-decoder can effectively transform the chart in image format into a CSV format (i.e. chart representation in Fig. 4). For intricate cognition tasks, we utilize LoRA [12] and fine-tune the pre-trained Vicuna-7B and Vicuna-13B as auxiliary decoders using text-text pair data including CSV, QA, summarization, and drawing codes.
Ultimately, two model variants are developed: ChartVLM-Base-7.3B (0.3B chart image encoder & base decoder + 7B auxiliary decoder) and ChartVLM-Large-8.3B (1.3B chart image encoder & base decoder + 7B auxiliary decoder). All the data we used during fine-tuning stage comes from ChartQA [27], PlotQA [29], Chart2Text [17], and SimChart9K [40]. Besides, the ChartVLM is trained using 32 $\times$ NVIDIA Tesla A100.
@@
LLaVA-1.5 [25] 13B 0.04 0.04 0.24 47.05 44.18 17.19 1.48 1.29 0.75
CogVLM [39] 18B 0.38 0.56 1.01 59.46 94.01 28.30 2.21 1.48 1.38
CogAgent [11] 10B 2.89 3.63 6.36 61.11 96.27 25.95 2.24 1.61 1.48
Monkey [20] 18B 0.00 0.00 0.00 66.84 94.44 21.61 1.85 1.62 1.24
QWen-VL [4] 9.6B 4.18 5.86 8.99 69.53 94.62 23.26 1.67 1.45 0.86
SPHINX-V2 [21] 13B 10.95 23.75 32.07 43.66 92.71 31.16 1.53 1.39 0.96
GPT-4V [31] - 20.91 26.00 36.09 70.43 95.22 33.04 3.17 3.12 2.63
@@
Deplot [22] 1.3B 8.89 19.04 24.08 - 89.84 - - - -
Matcha [23] 0.3B 0.92 1.10 1.16 5.03 7.90 14.41 - - -
ChartLlama [8] 13B 1.63 2.01 3.19 50.52 40.36 13.80 1.04 1.02 0.94
StructChart [40] 1.3B 0.46 0.94 1.77 - - - - - -
ChartAst [28] 13B 11.35 22.77 30.18 43.23 92.71 30.99 0.33 1.03 0.82
@@ 5.2 Baseline Models and Main Results
We select two kinds of MLLMs to make a comprehensive comparison. One group of MLLMs is made up of multi-modal large models, where models are trained towards general capability for various vision-language tasks. Here we select five of the most advanced MLLMs for evaluation comparison: LLaVA-1.5 [25], CogVLM [39], QWen-VL [4], SPHINX-V2 [21], and GPT-4V [31]. The other group of MLLMs represents the chart-related large models that are especially fine-tuned on chart-related tasks, including Deplot [22], Matcha [23], StructChart [40], ChartLlama [8], and ChartAssistant [28].
@@
Comparison with GPT-4V. As shown in Table 2, among all models, GPT-4V [31] is the only model that outperforms our ChartVLM in a few cognition tasks of the ChartX benchmark. This result is reasonable as GPT-4V is currently regarded as the most powerful MLLM for its strong ability to understand and describe information from images, e.g., summarization ability and description ability. However, for the perception tasks, since GPT-4V is a relatively general model, the structural extraction ability is inferior to our ChartVLM, which is specially designed for chart-related tasks. Furthermore, ChartVLM’s stronger ability to extract structural data from a chart image partially leads to a higher accuracy on the chart QA task (40.71%).
@@
Baseline VL-T5-OCR [27] - - 41.6
Tapas-OCR [27] - - 45.5
Pix2Struct [19] 81.6 30.5 56.0
MatCha [23] 90.2 38.2 64.2
Deplot [22] 69.3 36.6 52.9
ChartLlama [8] 90.4 49.0 69.7
GPT-4V [31] 76.1 64.5 70.3
@@
LLaVA-1.5 [25] 12.29 0.56 0.41 4.42 0.99 3.48 3.50 11.63 4.90
QWen-VL [4] 4.96 0.93 1.00 2.30 0.38 2.98 2.81 7.43 3.40
SPHINX-V2 [21] 5.53 1.51 1.21 2.75 1.38 3.96 4.09 9.73 4.79
Deplot [22] 3.82 - - 3.82 - - - - -
ChartLlama [8] 8.13 0.53 0.42 3.03 0.48 4.13 4.35 13.09 5.51
ChartAst [28] 55.24 3.55 1.37 20.05 3.81 6.06 6.04 34.14 12.51
@@
LLaVA-1.5 [25] 13B 0 0 0
QWen-VL [4] 9.6B 1.14 2.40 4.70
SPHINX-V2 [21] 13B 4.70 12.46 18.86
GPT-4V [31] - 14.35 19.00 27.22
Deplot [22] 1.3B 7.03 16.22 20.76
ChartLlama [8] 13B 1.39 1.68 2.37
ChartAst [28] 13B 5.99 14.93 21.19
@@
LLaVA-1.5 [25] 13B 0.04 0.04 0.24
QWen-VL [4] 9.6B 4.18 5.86 8.99
SPHINX-V2 [21] 13B 10.95 23.75 32.07
GPT-4V [31] - 20.91 26.00 36.09
Deplot [22] 1.3B 8.89 19.04 24.08
ChartLlama [8] 13B 1.63 2.01 3.19
ChartAst [28] 13B 11.35 22.77 30.18

### v6

@@ 1 Introduction
Versatile Multi-modal Large Language Models (MLLMs) (Bai et al., 2023; OpenAI, 2023; Achiam et al., 2023; Chen et al., 2024; Anthropic, 2024; Xia et al., 2024; Reid et al., 2024) have made promising progress in general-purpose vision-language applications such as multi-modal Question Answering (QA) (Lin et al., 2023; Bai et al., 2023; Wang et al., 2023; Liu et al., 2023b), embodied AI (Huang et al., 2023b), and mathematical reasoning (Trinh et al., 2024; Yang et al., 2023a; Jiang et al., 2022). Although MLLMs have demonstrated their powerful generalization ability in a wide range of multi-modal tasks, their performance in multi-modal reasoning tasks still falls short of human abilities (Yang et al., 2023b; Bubeck et al., 2023; Wang et al., 2024; Achiam et al., 2023). For instance, humans can easily extract numerical values from a given visual chart and engage in a series of complicated logical reasoning based on the extracted values. However, at present, the MLLMs’ ability to perform complicated logical reasoning based on chart data has not been fully explored.
@@
On top of this, the existing open-source chart datasets are consolidated for the training of ChartVLM, including ChartQA (Masry et al., 2022), Chart-to-text (Obeid and Hoque, 2020), PlotQA (Methani et al., 2020), and SimChart9K (Xia et al., 2023). Note that during the training process, ChartVLM has no access to any data from the ChartX evaluation set. Then, we conduct comprehensive comparisons of ChartVLM with current MLLMs (Bai et al., 2023; Liu et al., 2023b) on the ChartX evaluation set, including base abilities, e.g., data extraction, and advanced abilities, e.g., complicated problem-solving, where we demonstrate the superiority of our ChartVLM.
@@ 2 Related Work
Figure 1: Our work offers two insights: a) ChartX: a comprehensive multi-modal chart evaluation set encompassing 22 disciplinary topics, 18 chart types, and 7 tasks where models are evaluated using task-specific metrics such as EM, GPT-acc GPT-score, SCRM (Xia et al., 2023), and b) ChartVLM: a novel framework to perform the multi-tasks in the chart domain. Our key point is to leverage the instruction adapter to dynamically choose the task that needs to be executed. For downstream tasks that rely on querying chart information, we prioritize chart structural extraction before engaging in chart reasoning tasks. This sequence aims to enhance the interpretability of the reasoning results.
Table 1: Comparison with the existing chart-related benchmarks, where ChartX is constructed for comprehensively evaluating the off-the-shelf vision-language large models from more chart types and topics. Besides, EM denotes Exact Match and SCRM represents the Structuring Chart-oriented Representation Metric described in StructChart (Xia et al., 2023).
@@
PlotQA (Methani et al., 2020) N/A 3 1 33.7K PlotQA EM & AP
Chart-to-text (Obeid and Hoque, 2020) 6 6 1 6.6K Chart-to-text EM
ChartQA (Masry et al., 2022) 15 3 1 1.5K ChartQA EM
OpenCQA (Kantharaj et al., 2022) 10 5 1 1.2K OpenCQA EM & BLEU & ROUGE
@@
ChartLlama (Han et al., 2023) N/A 10 7 1.5K ChartQA & Chart-to-text EM & GPT
ChartBench (Xu et al., 2023) N/A 9 4 2K ChartBench Accuracy
MMC (Liu et al., 2023a) 5 6 9 2.1K MMC GPT
ChartAssisstant (Meng et al., 2024) N/A 9 5 1.5K ChartQA & OpenCQA EM& BLEU
@@
Chart Perception aims to extract the numerical and textual information from a given visual chart. By leveraging the OCR tools (Luo et al., 2021) to supplement the textual information, the basic function of extracting chart information can be achieved. Recently, some researchers (Hassan et al., 2023; Rane et al., 2021; Huang et al., 2023a) have attempted to perform a chart-to-table transformation for the visual chart perception task, by means of self-supervision from image-table pairs. For example, Deplot (Liu et al., 2022a) fine-tuned an image-to-text transformer for such conversion. StructChart (Xia et al., 2023) utilizes the encoder-decoder framework to achieve transformation. These methods extract the tabular format of a visual chart and leverage the external module such as GPT (Ouyang et al., 2022; Brown et al., 2020) to perform downstream tasks. However, their chart-related reasoning abilities strongly depend on external modules, whose scalability is hard to guarantee.
Chart Cognition is defined as a process to deal with intricate tasks related to both chart-related knowledge and common sense knowledge. A typical example is to query numerical points from a chart and give the prediction results using mathematical or logical reasoning. Recent studies (He et al., 2023; Tian et al., 2023; Zha et al., 2023; Lee et al., 2023; Liu et al., 2022b) focus on showing the reasoning ability of their models on chart domain. Pix2Struct (Lee et al., 2023) presents a pre-training method using masked screenshots from web pages, which is verified to be effective in chart understanding tasks such as ChartQA dataset (Masry et al., 2022). Besides, MatCha (Masry et al., 2022) decodes the answers to chart questions in an end-to-end manner, where the chart reasoning ability can be enhanced from MATH data (Saxton et al., 2019).
Multi-Modal Chart Generation and Benchmark. Chart data generation is a crucial step for scaling up the model ability (Tian et al., 2023; Liu et al., 2022b; Akhtar et al., 2023). Previous chart-related benchmarks only cover general three types of charts (line, pie, bar charts) and focus on a few tasks such as chart-to-table tasks for ChartQA (Masry et al., 2022), PlotQA (Methani et al., 2020), and Chart-to-Text (Obeid and Hoque, 2020), and QA tasks for DVQA (Kafle et al., 2018) and OpenCQA (Kantharaj et al., 2022). Recently, various benchmarks have been proposed in some works, e.g. MMC (Liu et al., 2023a), ChartLlama (Han et al., 2023), ChartBench (Xu et al., 2023), and ChartAssisstant (Meng et al., 2024), with the common characteristics of more types, more tasks, and more modalities of chart data, which is insightful for the chart community. However, as shown in Table 1, the data and metric diversity of charts used for evaluating multi-modal large models is relatively limited. For example, ChartBench (Xu et al., 2023) merely uses a two-sided judgment (yes or no) to evaluate model performance. The types of charts and data in MMC (Liu et al., 2023a) are also insufficient in verifying the chart ability of the off-the-shelf MLLMs.
@@ 3.1 Coverage Analysis of the Evaluation Set
Chart Tasks. Unlike previous chart benchmarks focusing on the category of visual logic reasoning tasks, the ChartX benchmark emphasizes the interpretability for all downstream chart-related tasks. Given that interpretability relies heavily on the ability to perceive chart information, ChartX categorizes perception-related tasks as base tasks, including title perception, chart type recognition, and Structural Extraction (SE). On the other hand, other chart-related tasks are classified as intricate cognition tasks, including chart-related Question Answering (QA), Chart Description, Chart Summarization, and Chart Redrawing. In the context of ChartX, QA refers to answering questions that are formulated solely based on the chart data, requiring reasoning derived directly from the provided chart information. This characteristic distinguishes ChartX from previous chart-related QA datasets like ChartQA (Masry et al., 2022). In ChartQA (Masry et al., 2022), there exists a certain number of QA pairs that cannot be answered solely based on the information presented in the given chart image. Chart Description aims at presenting detailed information and some insights from the distribution of chart data, while Chart Summarization features summarizing the trend-like or high-level characteristics from the given data in a few sentences. Chart Redrawing refers to plotting the given data into a new chart image with the same chart type of original data. The distribution of each task is listed in Fig. 1. For each image, together with labels of base tasks, we collect two QA samples, one description sample, one summarization sample, and one redrawing code sample. Overall, the samples from multi-tasks reach 48K in ChartX.
@@ 3.3 Two-stage Chart Data Generation
Utilizing the strong generation capabilities of GPT-4 (Achiam et al., 2023), ChartX is created through an automated online generation process with manual instructions. This involves a data-centric two-stage generation paradigm, encompassing the creation of perception and cognition data.
@@ 3.4 Task Evaluation Metrics
SCRM. Given that data in the chart has matrix-like row-column transformation invariance and transpose transformation invariance, Structuring Chart-oriented Representation Metric (SCRM) Xia et al. (2023) is employed to evaluate the extracted chart information (i.e. SE task), in which the linearized CSV tokens predicted by all models will be transformed to triplet format for performing SCRM evaluation.
@@ 4.3 Cascaded Decoders Design
For basic perception tasks, we fine-tune all the network weights pre-trained from Pix2Struct-base and Pix2Struct-large (Lee et al., 2023) model, using image-CSV pair data. The fine-tuned encoder and decoder are regarded as chart image encoder and base decoder in ChartVLM. After the fine-tuning stage is completed, the encoder-decoder can effectively transform the chart in image format into a CSV format (i.e. chart representation in Fig. 4). For intricate cognition tasks, we utilize LoRA (Hu et al., 2021) and fine-tune the pre-trained Vicuna-7B and Vicuna-13B as auxiliary decoders using text-text pair data including CSV, QA, summarization, and drawing codes.
Ultimately, two model variants are developed: ChartVLM-Base-7.3B (0.3B chart image encoder & base decoder + 7B auxiliary decoder) and ChartVLM-Large-8.3B (1.3B chart image encoder & base decoder + 7B auxiliary decoder). All the data we used during fine-tuning stage comes from ChartQA Masry et al. (2022), PlotQA Methani et al. (2020), Chart2Text Kanthara et al. (2022), and SimChart9K Xia et al. (2023). Besides, the ChartVLM is trained using 32 $\times$ NVIDIA Tesla A100.
@@
LLaVA-1.5 Liu et al. (2023b) 13B 0.04 0.04 0.24 47.05 44.18 17.19 1.48 1.29 0.75
CogVLM Wang et al. (2023) 18B 0.38 0.56 1.01 59.46 94.01 28.30 2.21 1.48 1.38
CogAgent Hong et al. (2023) 10B 2.89 3.63 6.36 61.11 96.27 25.95 2.24 1.61 1.48
Monkey Li et al. (2023) 18B 0.00 0.00 0.00 66.84 94.44 21.61 1.85 1.62 1.24
QWen-VL Bai et al. (2023) 9.6B 4.18 5.86 8.99 69.53 94.62 23.26 1.67 1.45 0.86
SPHINX-V2 Lin et al. (2023) 13B 10.95 23.75 32.07 43.66 92.71 31.16 1.53 1.39 0.96
GPT-4V OpenAI (2023) - 20.91 26.00 36.09 70.43 95.22 33.04 3.17 3.12 2.63
@@
Deplot Liu et al. (2022a) 1.3B 8.89 19.04 24.08 - 89.84 - - - -
Matcha Liu et al. (2022b) 0.3B 0.92 1.10 1.16 5.03 7.90 14.41 - - -
ChartLlama Han et al. (2023) 13B 1.63 2.01 3.19 50.52 40.36 13.80 1.04 1.02 0.94
StructChart Xia et al. (2023) 1.3B 0.46 0.94 1.77 - - - - - -
ChartAst Meng et al. (2024) 13B 11.35 22.77 30.18 43.23 92.71 30.99 0.33 1.03 0.82
@@ 5.2 Baseline Models and Main Results
We select two kinds of MLLMs to make a comprehensive comparison. One group of MLLMs is made up of multi-modal large models, where models are trained towards general capability for various vision-language tasks. Here we select five of the most advanced MLLMs for evaluation comparison: LLaVA-1.5 Liu et al. (2023b), CogVLM Wang et al. (2023), QWen-VL Bai et al. (2023), SPHINX-V2 Lin et al. (2023), and GPT-4V OpenAI (2023). The other group of MLLMs represents the chart-related large models that are especially fine-tuned on chart-related tasks, including Deplot Liu et al. (2022a), Matcha Liu et al. (2022b), StructChart Xia et al. (2023), ChartLlama Han et al. (2023), and ChartAssistant Meng et al. (2024).
@@
Comparison with GPT-4V. As shown in Table 2, among all models, GPT-4V (OpenAI, 2023) is the only model that outperforms our ChartVLM in a few cognition tasks of the ChartX benchmark. This result is reasonable as GPT-4V is currently regarded as the most powerful MLLM for its strong ability to understand and describe information from images, e.g., summarization ability and description ability. However, for the perception tasks, since GPT-4V is a relatively general model, the structural extraction ability is inferior to our ChartVLM, which is specially designed for chart-related tasks. Furthermore, ChartVLM’s stronger ability to extract structural data from a chart image partially leads to a higher accuracy on the chart QA task (40.71%).
@@
Baseline VL-T5-OCR Masry et al. (2022) - - 41.6
Tapas-OCR Masry et al. (2022) - - 45.5
Pix2Struct Lee et al. (2023) 81.6 30.5 56.0
MatCha Liu et al. (2022b) 90.2 38.2 64.2
Deplot Liu et al. (2022a) 69.3 36.6 52.9
ChartLlama Han et al. (2023) 90.4 49.0 69.7
GPT-4V OpenAI (2023) 76.1 64.5 70.3
@@
LLaVA-1.5 Liu et al. (2023b) 12.29 0.56 0.41 4.42 0.99 3.48 3.50 11.63 4.90
QWen-VL Bai et al. (2023) 4.96 0.93 1.00 2.30 0.38 2.98 2.81 7.43 3.40
SPHINX-V2 Lin et al. (2023) 5.53 1.51 1.21 2.75 1.38 3.96 4.09 9.73 4.79
Deplot Liu et al. (2022a) 3.82 - - 3.82 - - - - -
ChartLlama Han et al. (2023) 8.13 0.53 0.42 3.03 0.48 4.13 4.35 13.09 5.51
ChartAst Meng et al. (2024) 55.24 3.55 1.37 20.05 3.81 6.06 6.04 34.14 12.51
@@
LLaVA-1.5 Liu et al. (2023b) 13B 0 0 0
QWen-VL Bai et al. (2023) 9.6B 1.14 2.40 4.70
SPHINX-V2 Lin et al. (2023) 13B 4.70 12.46 18.86
GPT-4V OpenAI (2023) - 14.35 19.00 27.22
Deplot Liu et al. (2022a) 1.3B 7.03 16.22 20.76
ChartLlama Han et al. (2023) 13B 1.39 1.68 2.37
ChartAst Meng et al. (2024) 13B 5.99 14.93 21.19
@@
LLaVA-1.5 Liu et al. (2023b) 13B 0.04 0.04 0.24
QWen-VL Bai et al. (2023) 9.6B 4.18 5.86 8.99
SPHINX-V2 Lin et al. (2023) 13B 10.95 23.75 32.07
GPT-4V OpenAI (2023) - 20.91 26.00 36.09
Deplot Liu et al. (2022a) 1.3B 8.89 19.04 24.08
ChartLlama Han et al. (2023) 13B 1.63 2.01 3.19
ChartAst Meng et al. (2024) 13B 11.35 22.77 30.18